OpenStudy (anonymous):

What is the common difference in the arithmetic sequence 10, 8, 6, 4, ...?

OpenStudy (anonymous):

d = t2 - t1

OpenStudy (anonymous):

d = 8 - 10 = -2

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Well Done @Yahoo!

OpenStudy (anonymous):

Thanks